Find an equation of the line with the given points
(-5, - 7) (-3, -8)
I did
-8 - (-7)
-3 - (-5) = -1/2 (correct?)
And then I am not sure what to do to get it in the y = mx + b form
I believe it would be y = -1/2x +b But how do i get that b?
Thank you

Answer by Alan3354(69443)

Find an equation of the line with the given points
(-5,-7) (-3,-8)
I did
-8 - (-7)
-3 - (-5) = -1/2 (correct?)
And then I am not sure what to do to get it in the y = mx + b form
I believe it would be y = -1/2x +b But how do i get that b?
----------
Use y = (-1/2)x + b and either point to find b.